OpenStudy (anonymous):

Please help precal Find f(3) if f(x)=x^3+2x^2-x-1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

where you see an \(x\) replace it by \(3\)

OpenStudy (anonymous):

Ok thanks misty1212 I got 27

OpenStudy (anonymous):

59?

OpenStudy (anonymous):

i realy dont know the answer but if i did i would help u

TheSmartOne (thesmartone):

How did you get that?

OpenStudy (sleepyjess):

f(3) = 3^3 + 2(3)^3 - 3 - 1

TheSmartOne (thesmartone):

\(\sf\Large f(3)=(3)^2+2(3)^2-(3)-1\)

TheSmartOne (thesmartone):

Soooo what is \(\sf\Huge 3^3=?\) Hint: \(\sf\Huge 3^3 = 3 \times 3\times 3\)

OpenStudy (anonymous):

27

TheSmartOne (thesmartone):

and what is \(\sf\Huge 2\times 3^2=?\)

TheSmartOne (thesmartone):

This question isn't even pre-calc level tbh :P

OpenStudy (anonymous):

Ok thanks got it

TheSmartOne (thesmartone):

No problem. :)